OpenStudy (anonymous):

find the zeros of this function. state the multiplicity of multiple zeros y(3x+2)^3(x-5)^5

OpenStudy (anonymous):

So our function is: \(y(x)=(3x+2)^3(x-5)^5\). As we know, a zero of \(y(x)\) a value \(x\) such that \(y(x) = 0\). Since we've factored out our function into a product, at least one factor must be \(0\) for the product (i.e. the result of the function) to be \(0\); thus, to find the zeroes, we only need to solve for the \(x\)s that could result in one of the factors being \(0\).$$(3x+2)^3=0\\3x+2=0\\3x=-2\\x=-\frac23\\\ \\(x-5)^5=0\\x-5=0\\x=5$$The multiplicity of a zero \(x\) refers to the number of factors that using said \(x\) would cause to equal \(0\). Since \((3x+2)^3\) is really just \((3x+2)(3x+2)(3x+2)\), i.e. \((3x+2)\) multiplied with itself thrice, the zero \(-\frac23\) has multiplicity \(3\). Similarly, the zero \(5\) has multiplicity 5 (since \((x-5)^5=(x-5)(x-5)(x-5)(x-5)(x-5)\)).